DISTRIBUTION OF INTERNAL STATES OF CO AND OH FROM O(^{1}D) + C_{6}H_{6} AND C_{6}D_{6} DETERMINED WITH TIME-RESOLVED FOURIER-TRANSFORM SPECTROSCOPY

\underline{HUI-FEN CHEN}, National Tsing Hua University, Departmeny of Chemistry, Hsinchu 30013, Taiwan; and YUAN-PERN LEE, National Chiao Tung University, Department of Applied Chemistry, Hsinchu 30010, Taiwan, and Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan.

Rotationally resolved infrared emission spectra of CO (1 ≤ v ≤ 6) in the spectral region 1800 – 2350 cm^{−1} and OH (1 ≤ v ≤ 3) in the region 2800 – 3700 cm^{−1} were recorded with a step-scan Fourier-transform spectrometer. CO shows a rotational distribution corresponding approximately to temperatures 1520 K for v = 1 and 860 K for v = 2 – 6, with an average rotational energy of 9 ± 1 kJ mol^{−1} at the earliest applicable period (2.5 – 7.5 μs) upon photolysis. Extrapolation to time zero based on data in the range 2.5 – 27.5 μs yields an average nascent rotational energy of 14 ± 4 kJ mol^{−1}. Observed vibrational distribution of CO corresponds to a vibrational temperature of 5800 ± 330 K and an average vibrational energy of 33 ± 3 kJ mol^{−1}. OH shows a rotational distribution corresponding to temperatures 550 K for the P1 branch (v = 1 – 3) and 620 K for the P2 branch (v = 1 – 3), with an average nascent rotational energy of 4 ± 1 kJ mol^{−1}. The observed vibrational temperature of OH is 4830 ± 230 K, corresponding to an average vibrational energy of 21 ± 4 kJ mol^{−1}. The branching ratio of [CO]/[OH] is 2.1 ± 0.1 for O (^{1}D) + C_{6}H_{6} and no OD was observed from O (^{1}D) + C_{6}D_{6}. The significant deuterium isotope effect will be discussed.